A disk of radius 1.20cm. It has a charged of -6.0nC spread uniformly over the said disk. Find the magnitude and direction of the electric field this disk produces at point A on the axis of the disk a distance of 2.0cm from its center

1
Expert's answer
2022-05-18T08:32:10-0400
"E=k\frac{2q}{R^2}\left[1-\frac{z}{\sqrt{z^2+R^2}}\right]"

"E=9*10^9\frac{2*6.0*10^{-9}}{(0.012)^2}\left[1-\frac{0.02}{\sqrt{0.02^2+0.012^2}}\right]"

"E=5.34*10^4\:\rm N/C"